Find the domain of the function:
f(x)=(x)/(x^(2)+2x-35)
Answer by lwsshak3(11628) 
 Find the domain of the function:
f(x)=(x)/(x^(2)+2x-35)
**
To find the domain, set the denominator=0, then solve for x
x^2+2x-35=0
(x-7)(x+5)=0
x-7=0
x≠7
or
x+5=0
x≠-5
Domain: (-∞,-5) U (-5, 7) U (7,∞) (Interval notation)
{ x|x≠-5, ≠7} (set notation)
|
|
|